OpenStudy (anonymous):

which expression has the greatest value -4+8 or -2+(-3) or 1+2 or 4+(-6) plzzz need help for homework

OpenStudy (anonymous):

-4+8=4 -2+(-3)=-5 1+2=3 4+(-6)=-2 Thus -4+8 has the greatest value